OpenStudy (anonymous):

What is the slope of the line that contains the points (2, 6) and (-1, -6)?

jimthompson5910 (jim_thompson5910):

I'll get you started \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] \[\Large m = \frac{-6-6}{-1-2}\] \[\Large m = ???\]

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ 6}})\quad &({\color{red}{ -1}}\quad ,&{\color{blue}{ -6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

hold do you type color in Latex!!? D:

OpenStudy (anonymous):

\[\frac{ -12 }{ -3}\]

jimthompson5910 (jim_thompson5910):

reduce that to get ???

jimthompson5910 (jim_thompson5910):

what is 12/3 equal to?

OpenStudy (anonymous):

4!

jimthompson5910 (jim_thompson5910):

the two negatives cancel, so the result is positive therefore, the slope is 4

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome